Solve by the method of underminded coefficient using superposition apporoch

(D^2 +2D+2)y=2e^-x


1
Expert's answer
2021-10-15T11:43:54-0400

Homogeneous differential equation


"(D^2 +2D+2)y=0"

Corresponding (auxiliary) equation


"r^2+2r+2=0"

"(r+1)^2=-1"

"r=-1\pm i"

The general solution of the homogeneous differential equation is


"y_h=c_1e^{-x}\cos x+c_2e^{-x}\sin x"


Find the particular solution of the nonomogeneous differential equation


"y_p=Ae^{-x}"




"y_p'=-Ae^{-x}"

"y_p''=Ae^{-x}"

Substitute


"Ae^{-x}-2Ae^{-x}+2Ae^{-x}=2e^{-x}"

"A=2"

"y_p=2e^{-x}"

The general solution of the given nonhomogeneous differential equation is


"y=c_1e^{-x}\cos x+c_2e^{-x}\sin x+2e^{-x}"